A daredevil jumps a canyon 14 m wide. To do so, he drives a car up a 19° incline. What minimum speed must he achieve to clear the canyon? If the daredevil jumps at this minimum speed, what will his speed be when he reaches the other side?
1
Expert's answer
2012-10-09T09:23:36-0400
g*(t/2)=V0y=V0sin19 (That means, that after half of time we will reach maximum height

So, t=2V0sin19/g

S=Vox*t

V0x=v0*cos19

S=V0cos19*2v0sin19/g

S=(V0^2)sin38/g

V0=Sqrt(Sg/sin38)

V0=Sqrt(140/sin38)